OpenStudy (anonymous):

FIRST: Sal's Sandwich Shop sells wraps and sandwiches as part of its lunch specials. The profit on every sandwich is $2 and the profit on every wrap is $3. Sal made a profit of $1,470 from lunch specials last month. The equation 2x + 3y = 1,470 represents Sal's profits last month, where x is the number of sandwich lunch specials sold and y is the number of wrap lunch specials sold. So, there asking me to graph the function. On the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work or you may use graphing technology.

OpenStudy (anonymous):

I've worked out all the math: 2x+ 3y = 1470 3y= - 2x + 1470 y = -2/3 + 490 How would I graph this? Can someone show me please?!?! The intercept for is 490 and the slope is -2x/ 3